If X is a random variable with Poisson distribution satisfying P(X=0) = P(X=1), calculate E(X). 


1
Expert's answer
2021-06-14T15:45:52-0400

"P(X=x) = \frac{e^{-λ}λ^x}{x!} \\

x=0,1,2,… \\

E(x) = Var(x)= λ \\

P(X=0)=P(X=1) \\

\frac{e^{-λ}λ^0}{0!} = \frac{e^{-λ}λ^1}{1!} \\

λ=1 \\

E(x) = λ=1"